OpenStudy (anonymous):

method of undetermined coefficients 4y''+8y'+13y=1, y(0)=0, y'(0)=1/13

OpenStudy (anonymous):

use the quadratic formula find the roots to the characteristic polynomial \[4s ^{2}+8s+13=0\] then find all the solns to the corresponding homogenous D E . find a particular solution to the DE and put them together

OpenStudy (anonymous):

i got to the point \[\left( -8\pm \sqrt{-144} \right)\div8\] bit confused with the negative square root

OpenStudy (anonymous):

do i have to get imaginary numbers involved?

OpenStudy (anonymous):

\[\sqrt{-144}= 12i\]

OpenStudy (anonymous):

\[y _{p}=1/13\] is a particular soln to the DE

OpenStudy (anonymous):

the homo soln will be \[y _{h}(t)= c _{1}e ^{t}\cos8/12t+c _{2}e ^{t} \sin8/12\]

OpenStudy (anonymous):

add the part. soln to this, then use your init values to solve for \[c _{1}, c _{2}\]

OpenStudy (anonymous):

when u put 8/12 do you mean 12/8?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

and can you put it in brackets please?

OpenStudy (anonymous):

actually, i get that now, dont worry bout brackets

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for all the help, i have the final answer as \[y(x)=\exp(x)[(2/39)\sin ((3/2)x)]+(1/13)\] do you know if that is right?